OpenStudy (calculusxy):

MEDAL! A grocery store sells different types of trail mix. The cost and weight of each type is shown in the table. Which trail mix has the better price? Trail Mix A - $5 ( 3/4 lb) Trail Mix B - $7.50 ( 1 lb) Trail Mix C - $3.25 (4 oz)

OpenStudy (calculusxy):

@cwrw238

OpenStudy (calculusxy):

I got Trail Mix A.

OpenStudy (kendricklamar2014):

1 pound = 16 ounces

OpenStudy (kendricklamar2014):

3/4 pound = 12 ounces

OpenStudy (kendricklamar2014):

So ur right it is A

OpenStudy (calculusxy):

I have another question of the same type.

OpenStudy (kendricklamar2014):

Fire Away

OpenStudy (calculusxy):

Trail Mix A - $6 ( 3/4 lb) Trail Mix B - $8 (1 lb) Trail Mix B - $2.25 ( 1/4 lb)

OpenStudy (calculusxy):

I meant Trail Mix C for the last one.

OpenStudy (calculusxy):

I got Trail Mix A but my teacher says that it is supposed to be Trail Mix C.

OpenStudy (calculusxy):

I converted everything to unit rates.

OpenStudy (calculusxy):

I put: \[\frac{ $ }{ pound }\]

OpenStudy (calculusxy):

So to get one pound, I divided both the numerator and the denominator by 3/4 , 1 , 1/4.

OpenStudy (kendricklamar2014):

I think its A

OpenStudy (calculusxy):

\[\frac{ $6 }{ \frac{ 3 }{ 4 } } = \frac{ 6 }{ 1 } \times \frac{ 4 }{ 3 } = \frac{ 8 }{ 1 } = $8\]

OpenStudy (calculusxy):

Me too. But I don't know why my teacher says that it is C.

OpenStudy (kendricklamar2014):

me neither??